OpenStudy (mendicant_bias):

(ODE) I'm trying to solve a problem where I need to verify that the given functions form a fundamental set of solutions of the given DiffEq on the given interval, and form the general solution. Prompt below.

OpenStudy (mendicant_bias):

http://i.imgur.com/tbLzmgl.png

OpenStudy (mendicant_bias):

@ganeshie8 , is this where I just take the derivatives of the given set of solutions and plug them in to see if the equality holds true? Or, I dunno, how do I verify it otherwise? And how do I form the general solution?

ganeshie8 (ganeshie8):

yes the given solutions are independent so just verifying them will do

OpenStudy (mendicant_bias):

(By the way, this is before the introduction of the Wronskian matrix, so we can't use that/it's not relevant)

OpenStudy (unklerhaukus):

attach some constants to get the general solution

ganeshie8 (ganeshie8):

\[y = c_1e^{-3x} + c_2e^{4x}\]

OpenStudy (mendicant_bias):

^That sounds insultingly easy, lol, that can't be true.

ganeshie8 (ganeshie8):

question doesn't ask you to prove the independence, it just wants you to to check whethery they work or not

ganeshie8 (ganeshie8):

if they work, then you can simply take their linear combination for the general solution as above ^

OpenStudy (mendicant_bias):

Nevermind, this is during the section that introduces the Wronskian matrix, and the answer is phrased in the form of a Wronskian, you need to use it.

OpenStudy (mendicant_bias):

Okay, setting up the Wronskian: Just need to remember how to write matrices in LaTeX in this type of "editor".

OpenStudy (mendicant_bias):

God, I'm feeling like going to bed, but I need to get all this stuff done before. @ganeshie8 , do you know how I setup Matrices in OS LaTeX?

OpenStudy (mendicant_bias):

(Brb for 2 minutes, computer time about to expire like right now, renewing)

ganeshie8 (ganeshie8):

try this ``` \left| \begin{array}{} a&b\\ c&d\\ \end{array} \right| ```

ganeshie8 (ganeshie8):

that gives \[ \left| \begin{array}{} a&b\\ c&d\\ \end{array} \right| \]

OpenStudy (unklerhaukus):

``` \[ W(y_1,y_2)= \begin{vmatrix} y_1&y_2\\ y_1'&y_2' \end{vmatrix} \] ``` \[ W(y_1,y_2)= \begin{vmatrix} y_1&y_2\\ y_1'&y_2' \end{vmatrix} \]

OpenStudy (mendicant_bias):

Alright, renewed the computer, I'm back.

OpenStudy (mendicant_bias):

\[W(y_{1},y_{2})= \begin{vmatrix} e^{-3x}&e^{4x}\\ -3e^{-3x}&4e^{4x} \end{vmatrix} = 4e^x+3e^x\]

OpenStudy (mendicant_bias):

\[=7e^x\]

OpenStudy (mendicant_bias):

\[ 7e^{x} \neq 0, \]Solutions are linearly independent.

OpenStudy (mendicant_bias):

Alright, cool. New type of problem.